An object moves along a straight line path from (0,0) to the point (3,3) under the action of a force

f[vec{F} = (4 N/m)y hat{i} + (4 N/m)x hat{j}f]

Calculate the work done. (You must include all steps to get full credit for this problem.)
Answer #1

slope of the straight line = y2-y1-3-0 3-0 equation of straight line y-mr displacement vector r = xi + yj dy dr work done W =